OpenStudy (anonymous):

find the equation of the circle for which endpoints of the diameter are (-6, -3) and (4, 7)

OpenStudy (anonymous):

Find the midpoint use the coordinate with the distance-like formula to find the equation of the circle

OpenStudy (akashdeepdeb):

Use distance formula to get the length of the diameter: \[\sqrt{(x-x_1)^2 + (y-y_1)^2} = d\] \[d = 2r\] Find \(r\) Then use the two points to get the midpoint using midpoint formula:\[(x_{mid},y_{mid}) = (\frac{(x+x_1)}{2},\frac{(y+y_1)}{2})\] Thus, you now have your radius, and the midpoint co-ordinates. Use that in the circle equation: \[(x-x_{mid})^2 + (y-y_{mid})^2 = r^2\] Getting this? :)